# Operator Spreading and the Absence of Many-Body Localization

### v2

@@ III.1.1 No disorder
Figure 1: XX model: Without disorder, the number of distinct terms $n(k)$ in $[H,\sigma^{z}_{0}]^{(k)}$ grows as $(k+1)^{2}$ while it grows according to Eq. (III.13) with disorder. The general bound (II.18), valid for any free-fermion model, is also shown.
@@
$s^{(2)}(k)\equiv||[H,\sigma^{z}_{0}]^{(k)}||^{2}_{2}=2\sum_{i=1}^{k+1}\sum_{j=1}^{\frac{k+1}{2}}(C_{k})_{ij}^{2}=2^{2k}{{2k}\choose{k}}^{2}.$ (III.9)
@@
This means that all the terms contribute equally independent of the length of their support $l$ which clearly indicates delocalization. This analytical result can be checked by symbolic manipulations, see Fig. 3.
Figure 3: For $k\gg l$ the norm of operators with support on $l$ sites is asymptotically described by Eq. (III.12) and is independent of $l$.
@@
Figure 5: $s^{(1)}(k)$ and $s^{(2)}(k)$ for the Anderson model with $D=2$ and $D=10$. The results are averaged over $100$ samples. Note that the $2$-norm is only slightly smaller than the $1$-norm and the two norms are almost indistinguishable on this scale. The bound (III.14) is shown as well.
To obtain $s^{(1)}(k)$, we have to sum over all weighted walks through the graph which start at the top and have length $k$ by multiplying the values of all traversed edges, irrespective of the vertex the walk ends on. If, on the other hand, we want to calculate only the contributions of terms which have support on $l$ sites, $s^{(1)}_{l}(k)$, then we have to sum only over those weighted walks with $k$ steps which end at a vertex with terms on $l$ sites. We discuss this approach to calculating or bounding the norm of the commutator in more detail in App. B. Fig. 4 also already shows how the graph has to be modified in the Anderson case which we discuss next.
@@ III.1.2 Anderson Localization
What we can do exactly, however, is to count the number of distinct terms $n(k)$ in the commutator which occur at order $k$ and which is independent of the specific disorder realization. This is easiest done by making use of the graph Fig. 4 and by taking also into account the different possibilities where in the lattice these terms can be located. Doing so we find
$n(k)=2k(k+1)+\left\{\begin{array}[]{cc}1&\mbox{even}\\ 1-k&\mbox{odd.}\end{array}\right.$ (III.13)
This result is compared to symbolic calculations in Fig. 1. We stress once more that Eq. (III.13) is independent of the type of disorder or the specific disorder configuration and is thus also valid in the Aubry-André case discussed in Sec. III.1.3. Note that while the precise number of terms has changed as compared to the non-disordered case, importantly the number of terms still only increases quadratically with $k$.
Figure 6: Left column: Norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ for a fixed $k$ as a function of $l$ for the Anderson model with $D=2$ (top row) and $D=10$ (bottom row). Right column: $s^{(1)}_{l}(k)$ for a fixed $l$ as a function of $k$. The data are averaged over $100$ samples.
@@
$s^{(1)}(k)\leq 8(8+8D)^{k-1}=8^{k}(D+1)^{k-1}.$ (III.14)
As expected, we find that the norm has an exponential bound. Note that this bound increases with increasing disorder $D$, see Fig. 5. Also note that for $D=0$ we obtain the exact result for the norm in the non-disordered case, Eq. (III.8).
To see the difference between a free-fermionic model, where an initially localized operator spreads over the entire chain, and a system where the operator remains localized it is thus not sufficient to study the total operator norm. Instead, we have to investigate the spatial structure of the terms in the commutator at order $k$. Based on the graph in Fig. 4, we can derive a bound for the contribution of terms with support on $l$ sites to the norm at order $k$. We use that $|h_{j}|\leq D$. The derivation of this bound is discussed in detail in App. B.2. Here, we will consider the simplest possible approach where we only consider those terms with support on $l$ sites which have the largest possible number of factors $h_{j}$. This approach is justified for $D\gg 1$ and, as we will argue, gives the same asymptotics in the limit $k\gg l\gg 1$ as the more sophisticated approach discussed in App. B.2. Let us first consider terms which have support on a single site, i.e., which are of the form $\sigma^{z}_{j}$. As is obvious from the graph in Fig. 4, such terms are only generated at even orders of $k$. We collect the largest number of factors $h_{j}$ if we go down one level, then travel $k-2$ times along the vertical direction, and then go back up to the top of the graph. The $1$-norm of such terms is thus bounded by
$s^{(1)}_{1}(k)\leq 8\times(8D)^{k-2}\times 4=2^{2k+1}(2D)^{k-2}\,.$ (III.15)
@@
$\frac{s_{l}(k)}{s_{1}(k)}\lesssim\binom{k-1}{l-2}(2D)^{3-l}\approx\frac{1}{\sqrt{2\pi l}}\left(\frac{ke}{l}\right)^{l}e^{-\frac{l^{2}}{2k}}(2D)^{3-l}\,.$ (III.17)
We note that while for a fixed $k$ we indeed find that the ratio (III.17) goes exponentially to zero for large $l$, there is a maximum in the ratio which shifts to larger $l$ with increasing $k$. This is not quite the behavior we expect. Instead, we expect that with increasing $k$ the ratio (III.17) converges to a function which is exponentially decreasing in $l$. Here it is important to note that (III.17) is only an upper bound. In particular, we have only considered those paths which have a maximum number of $h_{j}$ factors. There are, however, many additional paths which can also interefere with each other and we expect that the prefactor in (III.17) will eventually converge with $k$ for a given $l$. This is consistent with what we observe from symbolic manipulations to finite orders of $k$, see Fig. 6. We note that in contrast to the non-disordered case where $s_{l}(k)$ grows exponentially and independent of $l$ for $k\gg l$, see Eq. (III.12) and Fig. 3, we find that with quenched disorder
$s_{l}(k)\sim\text{e}^{k}D^{-l}\,,$ (III.18)
@@ III.1.3 Aubry-André model
$h_{j}=\frac{\lambda}{2}\cos(2\pi\beta j+\phi)\,.$ (III.19)
Then, for $\beta$ Diophantine and almost any $\phi$—excluding a set of measure zero of resonant phases—the eigenstates of the Aubry-André model are localized for $\lambda>2$ [36, 37]. A common choice is to use the golden ratio $\beta=(1+\sqrt{5})/2$. The model thus provides an important test case where both a non-localized and a localized phase exist. Below the transition in the non-localized phase we expect that $s_{l}(k)$ is qualitatively described by Eq. (III.12), i.e., grows with $k$ independent of $l$ for $k\gg l$. In the localized phase, on the other hand, we expect that $s_{l}(k)$ is exponentially localized, see Eq. (III.18). Note that the graph for the Aubry-André model is the same as the one in the Anderson case shown in Fig. 4 with $D\to\lambda/2$ on the edges. With this replacement, also the bound (III.14) holds. We note that the considerations which let us to Eq. (III.18) in the Anderson case were valid only in the case of strong disorder. In this limit, they remain valid for the Aubry-André model. That the exponential decaying structure for the spatial support of the norm remains unchanged for all $D\neq 0$ in the Anderson case while there is a transition to the structure (III.12), which is independent of $l$ for $k\gg l$, for the Aubry-André model is beyond the strong disorder arguments we used in the previous section. All these results and expectations outlined above are fully consistent with the symbolic computations shown in Fig. 8.
Figure 8: Aubry-André model for $\beta=(1+\sqrt{5})/2$, $\phi=0$ where $\lambda=0.5$ (top row) and $\lambda=10$ (bottom row). For $\lambda=0.5$, $s^{(1)}_{l}(k)$ has the same exponential asymptotics independent of $l$ clearly showing that the model is not in the localized phase. In contrast, for $\lambda=10$ the norm scales as $s^{(1)}_{l}(k)\sim\text{e}^{k}\text{e}^{-l}$, i.e., is exponentially suppressed with increasing $l$ for $k$ fixed, demonstrating that the model is now in the localized phase.
@@ III.2.1 No disorder
Figure 9: In the interacting model with $\Delta=1$, the number of different terms $n(k)$ in the commutator $[H,\sigma_{0}^{z}]^{(k)}$ grows exponentially (symbols) and is very well fitted by $n(k)\sim 1.514\,k^{0.803}\exp(0.922k)$ in the non-disordered case and $n(k)\sim 1.492\,k^{0.815}\exp(0.998k)$ in the case with disorder (solid lines). The general bound (II.15) with $f=3$ and $f=4$ is shown as well.
The fit works extremely well with a least square deviation of $\chi^{2}\sim 10^{-2}$. Also shown in Fig. 9 is the general bound on $n(k)$, Eq. (II.15). The local operator basis for the Heisenberg model has dimension $f=4$ and consists of the operators ${\mathbbm{1},\sigma^{x},\sigma^{y},\sigma^{z}}$. We note, however, that only a small subset of terms in $[H,\sigma^{z}_{0}]^{(k)}$ contains an identity in between Pauli matrices. Ignoring such terms, we can set $f=3$ in Eq. (II.15) which gives a much tighter bound, see Fig. 9.
@@
Figure 11: The $2$-norm ratio $s^{(2)}(k+1)/s^{(2)}(k)$ for $D=0,5,12.5,25$. The data for $D=5$ and $D=12.5$ are averaged over $30$, and the data for $D=25$ over $50$ realizations. The norm ratio has to become a constant for large $k$ if the norm grows exponentially.
@@
Figure 12: The norm of operators $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ with support on $l$ sites for the Heisenberg model with disorder $D=0,5,12.5,25$. The data for $D=5$ and $D=12.5$ are averaged over $30$, the data for $D=25$ over $50$ disorder realizations. There is no indication for a convergence with increasing order $k$ to an exponential decay with $l$ even for very strong disorder. For $D=0$ and $D=25$ the results for the non-interacting case are shown for comparison (symbols with dotted lines).
Whereas in the non-interacting, non-disordered case the norm of the terms supported on $l$ sites grows asymptotically in the same way independent of $l$ if the order of the commutator $k$ is much larger than $l$, see Eq. (III.12), $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ has a maximum which shifts to larger and larger $l$ with increasing $k$ in the interacting case. For large $k$, the total norm is thus completely dominated by terms which have support on a large number of lattice sites. This is to be expected, given that the number of distinct terms in the commutator is increasing exponentially which also leads to a rapidly growing number of possible paths between those terms, see also the graph in Fig. 13.
@@
Figure 13: Graph for the $1$-norm $s^{(1)}(k)$ up to order $k=3$ for the Heisenberg model with random fields $h_{j}\in[-D,D]$. The number of terms grows exponentially with the length of their support.
@@ III.2.2 Many-body localization
From the perspective of operator spreading, localization in a many-body system is thus an extremely challenging requirement. In a graph with an equal number of edges at each level as realized in a non-interacting model, see Fig. 4, it is easy to see—at least in the limit of large disorder—how the norm of terms with support on only a few sites can dominate the total norm. The exponential growth of edges in the interacting case (see Fig. 13), on the other hand, suggests that the total norm for large $k$ will be entirely dominated by terms with support on many lattice sites. I.e., a local operator is generically expected to delocalize. This aspect also makes it difficult to see how a perturbative construction of local integrals of motion can be arranged. This is a point which seems to have not been sufficiently appreciated in Ref. [40, 41] where such a construction was suggested. We will get back to this point in Sec. IV. The only way a local operator can avoid to spread through the entire system in the many-body case thus seems to be intricate interference effects between the exponentially many different paths in the graph Fig. 13. We cannot prove that such interference effects do not exist. Instead, we will consider data from symbolic computations up to order $k=16$ and show that there is no evidence for localization. We note that at order $k$, we are considering operators which are spread over $12k+$ 1 sites in the lattice. I.e., at order $k=16$ we are dealing effectively with a chain of $33$ lattice sites which is well beyond what is possible in exact diagonalizations. We will also consider disorder strengths up to $D=25$ which is much larger than the critical values which have been suggested in the past [8, 10, 11, 42].
@@
Let us start by considering the number of distinct terms in the commutator at order $k$ which is shown in Fig. 9. We find that the number of terms grows faster in the disordered than in the case without disorder. This is easy to understand: additional terms appear in the commutator due to the random magnetic field terms while all the terms present before remain. The different classes of terms which appear up to order $k=3$ are shown in the graph in Fig. 13. Next, we consider the growth of the total norm which should be at most exponential in a localized phase as we have proven in Eq. (II.21), see also App. A. The data, however, are overall much better described by the general bound (II.12) than by an exponential, with the norm bound of the local Hamiltonian, $||h^{a}_{I}||\leq J$, used as a fitting parameter. While this does not exclude that for larger $k$ the scaling eventually does become exponential, we can say with confidence that no indications for localization are observed in spin chains up to $33$ sites. These system sizes are much larger than those used in exact diagonalizations, the method which claims of a many-body localized phase have been mostly based on [8, 10, 11, 19]. The stronger than exponential growth of the norm is further supported by considering the ratio of norms $s^{(2)}(k+1)/s^{(2)}(k)$ which is shown in Fig. 11. For the orders of the commutator which we can handle symbolically, there is no indication that this norm ratio saturates, even for extremely strong disorder, as it has to in a localized case where $s(k)\sim\exp(k)$. As the figure shows, the behavior in the interacting case is in stark contrast to the localized non-interacting case.
@@
Finally, let us also consider the norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ of terms with support on $l$ sites which is shown in Fig. 12 for various disorder strengths. Even for very strong disorder $D=25$ there is no indication of a convergence to an exponential decay with $l$. To the contrary, there is always a maximum which shifts to larger $l$ with increasing order $k$ just as in the case without disorder.
@@
Figure 14: Norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{3}(k)$ for a fixed $l$ as a function of $k$ for the Heisenberg model with $D=12.5$ (top) and $D=50$ (bottom). These results are to be compared with the non-interacting case shown in Fig. 7.

### v3

@@ III.1.1 No disorder
Figure 1: XX model: Without disorder, the number of distinct terms $n(k)$ in $[H,\sigma^{z}_{0}]^{(k)}$ grows as $(k+1)^{2}$ while it grows according to Eq. (III.14) with disorder. The general bound (II.18), valid for any free-fermion model, is also shown.
@@
$(s^{(2)}(k))^{2}\equiv||[H,\sigma^{z}_{0}]^{(k)}||^{2}_{2}=2\sum_{i=1}^{k+1}\sum_{j=1}^{\frac{k+1}{2}}(C_{k})_{ij}^{2}=2^{2k}{{2k}\choose{k}}^{2}.$ (III.9)
@@
This means that all the terms contribute equally independent of the length of their support $l$ which clearly indicates delocalization. We also note that the largest possible support at commutator order $k-1$ are $k$ sites with
$s_{k}^{(1)}(k-1)=2^{2k-1}\,.$ (III.13)
These analytical results can be checked by symbolic manipulations, see Fig. 3.
Figure 3: For $k\gg l$ the norm of operators with support on $l$ sites is asymptotically described by Eq. (III.12) and is independent of $l$. The exact value (III.13) when $s_{l}^{(1)}(k)$ first becomes non-zero is also shown.
@@
Figure 5: $s^{(1)}(k)$ and $s^{(2)}(k)$ for the Anderson model with $D=2$ and $D=10$. The results are averaged over $100$ samples. Note that the $2$-norm is only slightly smaller than the $1$-norm and the two norms are almost indistinguishable on this scale. The bound (III.15) is shown as well.
In this case, the graph is a $2$-regular graph with one node per level and $2$ edges per node. The constant number of nodes per level is directly reflected in the quadratic growth of the number of distinct terms $n(k)$.
To obtain $s^{(1)}(k)$, we have to sum over all weighted walks through the graph which start at the top and have length $k$ by multiplying the values of all traversed edges, irrespective of the vertex the walk ends on. If, on the other hand, we want to calculate only the contributions of terms which have support on $l$ sites, $s^{(1)}_{l}(k)$, then we have to sum only over those weighted walks with $k$ steps which end at a vertex with terms on $l$ sites. We discuss this approach to calculating or bounding the norm of the commutator in more detail in App. D. Fig. 4 also already shows how the graph has to be modified in the Anderson case which we discuss next.
@@ III.1.2 Anderson Localization
What we can do exactly, however, is to count the number of distinct terms $n(k)$ in the commutator which occur at order $k$ and which is independent of the specific disorder realization. This is easiest done by making use of the graph, see Fig. 4, and by taking also into account the different possibilities where in the lattice these terms can be located. Doing so we find
$n(k)=2k(k+1)+\left\{\begin{array}[]{cc}1&\mbox{even}\\ 1-k&\mbox{odd.}\end{array}\right.$ (III.14)
This result is compared to symbolic calculations in Fig. 1. We stress once more that Eq. (III.14) is independent of the type of disorder or the specific disorder configuration and is thus also valid in the Aubry-André case discussed in Sec. III.1.3. Note that while the precise number of terms has changed as compared to the non-disordered case, importantly the number of terms still only increases quadratically with $k$.
Figure 6: Left column: Norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ for a fixed $k$ as a function of $l$ for the Anderson model with $D=2$ (top row) and $D=10$ (bottom row). Right column: $s^{(1)}_{l}(k)$ for a fixed $l$ as a function of $k$. The data are averaged over $100$ samples. Note that the value of $s_{l}^{(1)}(k)$ when it first becomes non-zero at $k=l-1$ is not affected by disorder, see Eq. (III.13).
@@
$s^{(1)}(k)\leq 8(8+8D)^{k-1}=8^{k}(D+1)^{k-1}.$ (III.15)
As expected, we find that the norm has an exponential bound. Note that this bound increases with increasing disorder $D$, see Fig. 5. Also note that for $D=0$ we obtain the exact result for the norm in the non-disordered case, Eq. (III.8). Apart from the initial node and the node on the top right, the graph with disorder is a $3$-regular graph. I.e., each node is connected to three edges. We note that in every $n$-regular graph the norm $s^{(1)}(k)$ can not grow faster than exponential. This remains true even if the number of nodes increases with $k$ as, for example, in a Cayley tree as long as the number of edges per node is constant. This follows immediately from the fact that $s^{(1)}(k)$ is the sum of the product of edge values along all possible paths of length $k$. That $n(k)$ grows only quadratically means that the graph in the free-fermion case is particularly simple and has a constant number of nodes at each level.
To see the difference between a free-fermionic model, where an initially localized operator spreads over the entire chain, and a system where the operator remains localized it is thus not sufficient to study the total operator norm. Instead, we have to investigate the spatial structure of the terms in the commutator at order $k$. Based on the graph in Fig. 4, we can derive a bound for the contribution of terms with support on $l$ sites to the norm at order $k$. We use again that $|h_{j}|\leq D$. The derivation of this bound is discussed in detail in App. D.2. Here, we will consider the simplest possible approach where we only consider those terms with support on $l$ sites which have the largest possible number of factors $h_{j}$. This approach is justified for $D\gg 1$ and, as we will argue, gives the same asymptotics in the limit $k\gg l\gg 1$ as the more sophisticated approach discussed in App. D.2. Let us first consider terms which have support on a single site, i.e., which are of the form $\sigma^{z}_{j}$. As is obvious from the graph in Fig. 4, such terms are only generated at even orders of $k$. We collect the largest number of factors $h_{j}$ if we go down one level, then travel $k-2$ times along the vertical direction, and then go back up to the top of the graph. The $1$-norm of such terms is thus bounded by
$s^{(1)}_{1}(k)\leq 8\times(8D)^{k-2}\times 4=2^{2k+1}(2D)^{k-2}\,.$ (III.16)
@@
$\frac{s_{l}(k)}{s_{1}(k)}\lesssim\binom{k-1}{l-2}(2D)^{3-l}\approx\frac{1}{\sqrt{2\pi l}}\left(\frac{ke}{l}\right)^{l}e^{-\frac{l^{2}}{2k}}(2D)^{3-l}\,.$ (III.18)
We note that while for a fixed $k$ we indeed find that the ratio (III.18) goes exponentially to zero for large $l$, there is a maximum in the ratio which shifts to larger $l$ with increasing $k$. This is not quite the behavior we expect. Instead, we expect that with increasing $k$ the ratio (III.18) converges to a function which is exponentially decreasing in $l$. Here it is important to note that (III.18) is only an upper bound. In particular, we have only considered those paths which have a maximum number of $h_{j}$ factors. There are, however, many additional paths and we expect that the prefactor in (III.18) will eventually converge with $k$ for a given $l$. This is consistent with what we observe from symbolic manipulations to finite orders of $k$, see Fig. 6. We note that in contrast to the non-disordered case where $s_{l}(k)$ grows exponentially and independent of $l$ for $k\gg l$, see Eq. (III.12) and Fig. 3, we find that with quenched disorder
$s_{l}(k)\sim\text{e}^{k}D^{-l}\,,$ (III.19)
@@ III.1.3 Aubry-André model
$h_{j}=\frac{\lambda}{2}\cos(2\pi\beta j+\phi)\,.$ (III.20)
Then, for $\beta$ Diophantine and almost any $\phi$—excluding a set of measure zero of resonant phases—the eigenstates of the Aubry-André model are localized for $\lambda>2$ [43, 44]. A common choice is to use the golden ratio $\beta=(1+\sqrt{5})/2$. The model thus provides an important test case where both a non-localized and a localized phase exist. Below the transition in the non-localized phase we expect that $s_{l}(k)$ is qualitatively described by Eq. (III.12), i.e., grows with $k$ independent of $l$ for $k\gg l$. In the localized phase, on the other hand, we expect that $s_{l}(k)$ is exponentially localized, see Eq. (III.19). Note that the graph for the Aubry-André model is the same as the one in the Anderson case shown in Fig. 4 with $D\to\lambda/2$ on the edges. With this replacement, also the bound (III.15) holds. We note that the considerations which let us to Eq. (III.19) in the Anderson case were valid only in the case of strong disorder. In this limit, they remain valid for the Aubry-André model. That the exponential decaying structure for the spatial support of the norm remains unchanged for all $D\neq 0$ in the Anderson case while there is a transition to the structure (III.12), which is independent of $l$ for $k\gg l$, for the Aubry-André model is beyond the strong disorder arguments we used in the previous section. All these results and expectations outlined above are fully consistent with the symbolic computations shown in Fig. 8. Note that the data in the left panels for the non-localized phase are shown on a linear scale while the results in the right panels for the localized phase are shown on a logarithmic scale.
Figure 8: Aubry-André model for $\beta=(1+\sqrt{5})/2$, $\phi=0$ where $\lambda=0.5$ (left column, linear scale) and $\lambda=10$ (right column, logarithmic scale). For $\lambda=0.5$, $s^{(1)}_{l}(k)$ is extrapolated (blue lines) and the results are consistent with $s^{(1)}_{l}(k)\sim C\text{e}^{k}$ with a constant $C$ independent of $l$. This leads to $s^{(1)}_{l}(k\to\infty)/s^{(1)}_{3}(k\to\infty)\approx 1$ (bottom left panel) clearly showing that the model is not in the localized phase. In contrast, for $\lambda=10$ the norm scales as $s^{(1)}_{l}(k)\sim\text{e}^{k}\text{e}^{-l}$, i.e., is exponentially suppressed with increasing $l$ for $k$ fixed, demonstrating that the model is now in the localized phase (bottom right panel).
@@ III.2.1 No disorder
Figure 9: In the interacting model with $\Delta=1$, the number of different terms $n(k)$ in the commutator $[H,\sigma_{0}^{z}]^{(k)}$ grows exponentially (symbols) and is very well fitted by $n(k)\sim 1.514\,k^{0.803}\exp(0.922k)$ in the non-disordered case and $n(k)\sim 1.492\,k^{0.815}\exp(0.998k)$ in the case with disorder (solid lines). The general bounds (II.15) with $f=3$ and $f=4$ are shown as well.
The fit works extremely well with a least square deviation of $\chi^{2}\sim 10^{-2}$. The local operator basis for the Heisenberg model has dimension $f=4$ and consists of the operators ${\mathbbm{1},\sigma^{x},\sigma^{y},\sigma^{z}}$. We note, however, that only a small subset of terms in $[H,\sigma^{z}_{0}]^{(k)}$ contains an identity in between Pauli matrices. Ignoring such terms, we can set $f=3$ in Eq. (II.15) which gives a much tighter bound, see Fig. 9. The exponential growth of $n(k)$, instead of the quadratic growth in the free-fermion case, indicates that in the corresponding commutator graph the number of nodes per level is no longer constant.
@@
Figure 11: Average of the $2$-norm ratio $s^{(2)}(k)/s^{(2)}(k-1)$ for the XXZ chain with $\Delta=0$ and $\Delta=1$ for disorder strengths $D=0,5,12.5,25$. The data for $D=5$ and $D=12.5$ are averaged over $50$, and the data for $D=25$ over $70$ realizations. The shaded regions indicate the intervals into which 50% of all realizations fall. The norm ratio has to become a constant for large $k$ if the norm grows exponentially.
@@
Figure 12: The norm of operators $s^{(1)}_{l}(k)/E[s^{(1)}_{1}(k)]$ with support on $l$ sites for the Heisenberg model with disorder $D=0,5,12.5,25$. The data for $D=5$ and $D=12.5$ are averaged over $50$, the data for $D=25$ over $70$ disorder realizations. There is no indication for a convergence with increasing order $k$ to an exponential decay with $l$ even for very strong disorder. For $D=0$ and $D=25$ the results for the non-interacting case are shown for comparison (symbols with dotted lines).
Whereas in the non-interacting, non-disordered case the norm of the terms supported on $l$ sites grows asymptotically in the same way independent of $l$ if the order of the commutator $k$ is much larger than $l$, see Eq. (III.12), $s^{(1)}_{l}(k)/s^{(1)}_{1}(k)$ has a maximum which shifts to larger and larger $l$ with increasing $k$ in the interacting case. For large $k$, the total norm is thus completely dominated by terms which have support on a large number of lattice sites. This is to be expected, given that the number of distinct terms in the commutator is increasing exponentially which also leads to a rapidly growing number of possible paths between those terms, see also the graph in Fig. 13. Note that the graph is no longer regular; the number of edges per node is not constant and is increasing with the level the node is situated on. Note that this increase of the number of edges is a prerequisite for a faster than exponential growth of the norm.
@@
Figure 13: Graph for the $1$-norm $s^{(1)}(k)$ up to order $k=3$ for the Heisenberg model with random fields $h_{j}\in[-D,D]$. The number of nodes grows exponentially with the length of the support.
@@ III.2.2 Many-body localization
From the perspective of operator spreading, localization in a many-body system is thus an extremely challenging requirement. In a graph with an equal number of edges at each level as realized in a non-interacting model, see Fig. 4, it is easy to see—at least in the limit of large disorder—how the norm of terms with support on only a few sites can dominate the total norm. The exponential growth of edges in the interacting case (see Fig. 13), on the other hand, suggests that the total norm for large $k$ will be entirely dominated by terms with support on many lattice sites. I.e., a local operator is generically expected to delocalize. This aspect also makes it difficult to see how a perturbative construction of local integrals of motion can be arranged. This is a point which seems to have not been sufficiently appreciated in Refs. [47, 48] where such a construction was suggested. We will get back to this point in Sec. IV. The only way a local operator can avoid to spread through the entire system in the many-body case thus seems to be intricate interference effects between the exponentially many different paths in the graph, Fig. 13. While such fine-tuned destructive interference seems implausible as a generic mechanism for many-body localization we cannot, in general, show that such effects do not exist. We note, however, that for the transverse Ising chain with random fields—the model considered in the proof of MBL in Refs. [47, 48]—a gauge can be defined such that the commutator has a stoquastic-type property [37]. I.e., all terms in the commutator in this gauge have a positive amplitude. For this model, interference effects as a mechanism for localization are thus strictly excluded. For the Heisenberg model considered here, such a gauge does not exist. We will instead consider data from symbolic computations up to order $k=16$ and show that there is no evidence for destructive interference effects and a corresponding localization. We note that at order $k$, we are considering operators which are spread over $12k+$ 1 sites in the lattice. I.e., at order $k=16$ we are dealing effectively with a chain of $33$ lattice sites which is well beyond what is possible in exact diagonalizations. We will also consider disorder strengths up to $D=25$ which is much larger than the critical values which have been suggested in the past [8, 10, 11, 49].
@@
Let us start by considering the number of distinct terms in the commutator at order $k$ which is shown in Fig. 9. We find that the number of terms grows faster in the disordered than in the case without disorder. This is easy to understand: additional terms appear in the commutator due to the random magnetic field terms while all the terms present before remain. The different classes of terms which appear up to order $k=3$ are shown in the graph in Fig. 13. Next, we consider the growth of the total norm which should be at most exponential in a localized phase as we have proven in Eq. (II.21), see also App. A. The data, however, are overall much better described by the general bound (II.12) than by an exponential, with the norm bound of the local Hamiltonian, $||h^{a}_{I}||\leq J$, used as a fitting parameter. While this does not exclude the possibility that for larger $k$ the scaling eventually does become exponential, we can say with confidence that no indications for localization are observed in spin chains up to $33$ sites. These system sizes are much larger than those used in exact diagonalizations, the method which claims of a many-body localized phase have been mostly based on [8, 10, 11, 19]. The stronger than exponential growth of the norm is further supported by considering the ratio of norms $s^{(2)}(k)/s^{(2)}(k-1)$ which is shown in Fig. 11. For the orders of the commutator which we can handle symbolically, there is no indication that this norm ratio saturates, even for extremely strong disorder, as it has to in a localized case where $s(k)\sim\exp(k)$. As the figure shows, the behavior in the interacting case is in stark contrast to the localized non-interacting case.
@@
Finally, let us also consider the norm ratio $s^{(1)}_{l}(k)/E[s^{(1)}_{1}(k)]$ of terms with support on $l$ sites which is shown in Fig. 12 for various disorder strengths. Here $E[s^{(1)}_{1}(k)]$ denotes the mean of terms with support on one site. Even for very strong disorder $D=25$ there is no indication of a convergence to an exponential decay with $l$. To the contrary, there is always a maximum which shifts to larger $l$ with increasing order $k$ just as in the case without disorder.
@@
Figure 14: Norm ratio $s^{(1)}_{l}(k)/s^{(1)}_{3}(k)$ for a fixed $l$ as a function of $k$ for the Heisenberg model with $D=12.5$ (top) and $D=25$ (bottom). These results are to be compared with the non-interacting case shown in Fig. 7.